# Adv-Diffusion: Imperceptible Adversarial Face Identity Attack via Latent Diffusion Model

### 18 Dec 2023 v1

@@ 6 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62306227, Grant 62276198, Grant U22A2035, Grant U22A2096, Grant 62036007, Grant 61922066, Grant 61906143, Grant 61902297, Grant 61806152 and Grant 61876142; in part by the Key Research and Development Program of Shaanxi (Program No. 2023-YBGY-231); in part by Young Elite Scientists Sponsorship Program by CAST under Grant 2022QNRC001; in part by the Guangxi Natural Science Foundation Program under Grant 2021GXNSFDA075011; in part by Natural Science Basic Research Plan in Shaanxi Province of China under Grant 2022JQ-696; in part by the Technology Innovation Leading Program of Shaanxi under Grant 2022QFY01-15; in part by CCF-Tencent Open Fund, in part by Open Research Projects of Zhejiang Lab under Grant 2021KG0AB01, and in part by the Fundamental Research Funds for the Central Universities under Grant QTZX23083, and Grant XJS221502.
@@ Appendix B More Analysis of Semantic Appearance Deformation Degree
Appendix C Visualization Comparison and Analysis
@@
Appendix D Expanded Application Scope
Our method can be applied to non-face images for classification tasks conveniently. We implement the Adv-Diffusion method for images from the Internet. Here we use ResNet50 as the target model and directly input source images without masks to the latent encoder to get the condition c. As shown in the following figure, the generated adversarial samples contain both good attacking transferability and stealthiness when evaluated with ResNet101.
Appendix E Real World Application
We implement the proposed method for face images in the wild with large pose angles and different backgrounds as shown in the following figure. Experimental results prove the proposed method can be suitable in most real-world scenarios. For extremely complex scenes, we can leverage more advanced parsing models (or manually tuning) to improve the generation quality. Besides, the extended method can alleviate the effect of parsing performance, further proving the potential attacking performance in real applications.
Appendix F Negative Results
We add some negative results as shown in the following figure. We find that complex hair styling is more likely to bring adverse effects of the parsing and diffusion generative models. Thus, we will explore a more robust adversarial generative pipeline in future work.
Appendix G User Study

### 28 Dec 2023 v2

@@ 6 Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62306227, Grant 62276198, Grant U22A2035, Grant U22A2096, Grant 62036007, Grant 61922066, Grant 61906143, Grant 61902297, Grant 61806152 and Grant 61876142; in part by the Key Research and Development Program of Shaanxi (Program No. 2023-YBGY-231); in part by Young Elite Scientists Sponsorship Program by CAST under Grant 2022QNRC001; in part by the Guangxi Natural Science Foundation Program under Grant 2021GXNSFDA075011; in part by Natural Science Basic Research Plan in Shaanxi Province of China under Grant 2022JQ-696; in part by the Technology Innovation Leading Program of Shaanxi under Grant 2022QFY01-15; in part by CCF-Tencent Open Fund, in part by Open Research Projects of Zhejiang Lab under Grant 2021KG0AB01, and in part by the Fundamental Research Funds for the Central Universities under Grant QTZX23083, Grant XJS221502 and Grant QTZX23042.
@@ Appendix A Additional Experimental Details
Computation Burden.
The computational burden of the method is mainly related to the inference process. We conduct experiments on RTX 3090 GPU with a batch size of $2$, and each batch consumes about $62.1s$.
@@ Appendix B More Analysis of Semantic Appearance Deformation Degree
Appendix C Additional Experimental Analysis
Non-face image application
Our method can be applied to non-face images for classification tasks conveniently. We implement the Adv-Diffusion method for images from the Internet. Here we use ResNet50 as the target model and directly input source images without masks to the latent encoder to get the condition c. As shown in the following figure 2, the generated adversarial samples contain both good attacking transferability and stealthiness when evaluated with ResNet101.
Figure 2: The non-face image results of the proposed adversarial face generation algorithm.
Face generation in the wild
We implement the proposed method for face images in the wild with large pose angles and different backgrounds as shown in the following figure 3. Experimental results prove the proposed method can be suitable in most real-world scenarios. For extremely complex scenes, we can leverage more advanced parsing models (or manually tuning) to improve the generation quality. Besides, the extended method can alleviate the effect of parsing performance, further proving the potential attacking performance in real applications.
Figure 3: The face in the wild results of the proposed adversarial face generation algorithm.
Failure case
We add some negative results as shown in the following figure 4. We find that complex hair styling is more likely to bring adverse effects of the parsing and diffusion generative models. Thus, we will explore a more robust adversarial generative pipeline in future work.
Figure 4: The failure generation results of the proposed adversarial face generation algorithm.
User study
Appendix D Visualization Comparison and Analysis